OpenStudy (anonymous):

\[f(x,y) = 2^{x^2+y^2}\] draw level curves

OpenStudy (anonymous):

Usually I set k=f(x,y), solve for y, and then draw curves for different values of k What should I do in this case?

OpenStudy (anonymous):

\[k=2^{-(x^2+y^2)}\\log_2k=-(x^2+y^2)\\-log_2k=x^2+y^2\] Looks like they are going to be circles, but with a negative radius?

OpenStudy (anonymous):

x^2+y^2=ln(r^ln2)

OpenStudy (anonymous):

r=1 x^2+y^2=1 circle

OpenStudy (anonymous):



OpenStudy (anonymous):

r=2^ln2=2^0.69..=1.6168..

OpenStudy (anonymous):

how did you get \[\ln(r^{\ln2})\]

OpenStudy (anonymous):